OpenStudy (anonymous):

HELP!! MEDAL! Solve by completing the square. t2 - 8t - 5 = 0

OpenStudy (anonymous):

t² - 8t - 5 = 0 t² - 8t = 5 Now to complete the square: t² - 8t + (-8/2)² = 5 + (-8/2)² t² - 8t + 16 = 5 + 16 (t - 4)² = 21 Take the square root of both sides to find t

OpenStudy (anonymous):

did you got it ?